Find the equilibrium point for the demand and supply functions D9p)=64-4p and S(p)=32 + 4p.
Answer by stanbon(75887)

Find the equilibrium point for the demand and supply functions
D(p)=64-4p and S(p)=32 + 4p.
---
Solve D(p) = S(p)
64-4p = 32 + 4p
----
8p = 32
---
p = 4
==========
Cheers,
Stan H.